Please show work! Thank youQuestion 3 --/2 View Policies Current Attempt in Progress Wildhorse Company purchased equipment for $266,000 on October 1, 20
Answer #1

A. Straight Line Method

>> Depreciable asset = Cost of Asset - Salvage value = $ 266,000 - $ 14,000 = $ 252,000.

>> Depreciation per annum = Depreciable asset / Useful life of asset = $ 252,000 / 8 = $ 31,500.

>> Depreciation for 2020 = $ 31,500 * ( 3 / 12 ) = $ 7,875.

==========================================================================================

B. Activity Method - units produced

>> Depreciable asset = Cost of Asset - Salvage value = $ 266,000 - $ 14,000 = $ 252,000.

>> Depreciation per unit = Deprecible asset / total units = $ 252,000 / 40,000 = $ 6.3

>> Depreciation for 2020 = 1100 * $ 6.3 = $ 6,930.

==========================================================================================

C. Activity Method - Working hours

>> Depreciable asset = Cost of Asset - Salvage value = $ 266,000 - $ 14,000 = $ 252,000.

>> Depreciation per hour = Deprecible asset / total hours = $ 252,000 / 20,000 = $ 12.6

>> Depreciation for 2020 = 530 * $ 12.6 = $ 6,678.

==========================================================================================

D. Sum of years digit method

>> Depreciable asset = Cost of Asset - Salvage value = $ 266,000 - $ 14,000 = $ 252,000.

>> Sum of years = n ( n + 1 ) / 2 = 8 ( 8 + 1 ) / 2 = 36.

>> Depreciation for 2022 = $ 252,000 * ( 6.75 / 36 ) = $ 47,250.

==========================================================================================

E. Double declining method.

Depreciation rate = 200 / useful life = 200 / 8 = 25 %.

Depreciation for 2020 = $ 266,000 * 25 % * ( 3 /12 ) = $ 16,625.

Depreciation for 2021 = ( $ 266,000 - $ 16,625 ) * 25 % = $ 62,344